### v4

# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF [1] and cosine similarity [2] to identify related episodes and enhance the overall story structure. Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
Index Terms:
Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
**footnotetext: Equal contribution.
I Introduction
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [3, 4, 5, 6, 7, 8]. Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [9]. However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10]. For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
Similarly, [12] highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents. [13] conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks. Additionally, [14] introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society. These researches inspired the design of our new framework.
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) [15], which dynamically incorporates relevant context to enhance narrative coherence. Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
  • We introduce SCORE, an LLM-based evaluation framework that detects narrative inconsistencies in AI-generated stories.
  • We incorporate a Retrieval-Augmented Generation (RAG) approach, utilizing episode-level summaries and key item tracking to improve narrative coherence.
  • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
  • We outperform baseline GPT model [16] in detecting continuity errors and maintaining overall narrative coherence.
II Method
Our proposed method, SCORE, consists of three main components: (1) an LLM-based evaluation framework to assess the coherence of key story elements, (2) automatic generation of episode summaries to track plot development, and (3) a retrieval-augmented generation (RAG) approach that integrates the first two components, enabling enhanced user interaction and ensuring narrative consistency.
As the framework is intended solely for academic research purposes, its use is consistent with the original access conditions of all incorporated tools and data sources.
Figure 1: The SCORE framework for improving AI-generated story coherence. (a) Extracts key item statuses per episode. (b) Conducts detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
II-A Continuity Analysis and Key Item Status Correction
By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without narrative explanation. Let $S_{i}(t)$ represent the state of item $i$ at time $t$, where $S_{i}(t)\in\{\text{active},\text{lost},\text{destroyed}\}$. If item $i$ reappears at time $t_{k}$ with $S_{i}(t_{k})=\text{active}$ after being previously marked as $S_{i}(t_{k-1})\in\{\text{lost},\text{destroyed}\}$, we flag this as a continuity error. To maintain consistency, the state remains $S_{i}(t_{k-1})$, avoiding an incorrect update. This approach systematically corrects discrepancies in item states, ensuring that narrative continuity is preserved by preventing erroneous state transitions.
II-B Key Item Interaction Analysis
For each episode, we conduct a thorough evaluation by summarizing key plot points, character actions, and tracking interactions with important items. Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote interactions with key item $i$. The model generates summaries that encapsulate essential elements, including $A_{c}(t)$ (character actions), relationships, and emotional changes across the episode. It then analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis. This step aggregates relevant narrative information—combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$—to facilitate more precise future retrieval. The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
II-C Similarity-Based Episode Evaluation and Sentiment Analysis
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured JSON files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS [17] and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
Sentiment analysis is then applied to both the current and retrieved episodes. A sentiment score $\sigma(e)$, ranging from 0 to 1, is assigned to each episode $e$ by GPT-4, reflecting its emotional tone. These scores help refine the selection by ensuring both text similarity and sentiment consistency are considered, thus preventing errors from large sentiment discrepancies.
Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation. The focus is on narrative aspects such as character consistency, plot progression, emotional authenticity, and key item continuity. This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
TABLE I: Performance comparison of models with and without SCORE
Model Consistency Coherence Item Status Complex Question
GPT-4o-mini 78.2 76.7 0 24.56
+ SCORE 82.6 (↑4.4) 77.5 (↑0.8) 80.5 (↑80.5) 63.0 (↑38.44)
GPT-4o 86.78 82.21 0 76.32
+ SCORE 88.68 (↑1.9) 89.91 (↑7.7) 96 (↑96) 88.75 (↑12)
GPT-4 83.21 84.32 0 82.34
+ SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
III Experiments
To evaluate the effectiveness of SCORE, we conducted experiments on stories generated by large language models (LLMs). These experiments assessed the framework’s ability to maintain narrative coherence, detect continuity errors, and ensure emotional consistency throughout episodic storytelling.
III-A Dataset Preparation
We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy. Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements. This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
III-B Baselines
We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini. In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms. We used the same LLM-generated stories to evaluate different models. For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
III-C Evaluation Process
For each episode in the dataset, the evaluation process involved several key steps. First, we conducted an initial evaluation by directly uploading the files to ChatGPT, testing if GPT could assess the story correctly without any prompts. We used GPT-4o-mini, GPT-4o, and GPT-4 for narrative evaluation.
Next, we configured the LLM for a more detailed evaluation, using our preprocessed files to ensure accurate tracking of key items across episodes. We also employed the Retrieval-Augmented Generation (RAG) framework to retrieve episodes similar to the one being evaluated. By using FAISS to calculate cosine similarity scores between episodes, the system was able to retrieve episodes that were semantically and emotionally aligned with the current one, providing additional context. This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
III-D Metrics
We evaluated the framework based on several key metrics. Narrative coherence was assessed by examining how well the framework maintains the logic of the story. Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected. Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
III-E Results
Our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies. Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression. The retrieval-augmented generation process helped GPT better filter irrelevant information, understand the current story context, and improved its ability to detect narrative continuity across multiple episodes. When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
IV Limitations
While our SCORE framework offers significant improvements, it also presents several limitations:
  • Dependence on Retrieval Accuracy: The effectiveness of SCORE relies on accurately retrieving relevant episodes. Errors in similarity calculation or sentiment analysis may lead to the exclusion of important contextual information, affecting the overall evaluation quality.
  • Challenges in Capturing Emotional Nuances: Sentiment analysis may not fully capture the complexity of emotions, potentially missing subtle shifts or deeper emotional layers within the narrative.
  • Resource Intensive: Despite optimizations, SCORE still requires considerable computational resources for similarity calculations and sentiment analysis, which can be limiting for large datasets.
  • Generalizability: SCORE has been primarily tested on GPT-4-generated stories, and its performance on narratives from other models or with different structures remains an open question.
V Conclusion
We introduced SCORE, a novel LLM-based evaluation framework aimed at improving the coherence and emotional consistency of AI-generated stories. By incorporating Retrieval-Augmented Generation (RAG), TF-IDF, cosine similarity, and sentiment analysis, our method retrieves contextually relevant information, enabling more accurate and detailed assessments.
Experimental results demonstrate that SCORE outperforms traditional methods by effectively identifying continuity errors, maintaining narrative coherence, and providing valuable insights for episodic storytelling. However, limitations such as dependency on retrieval accuracy and computational demands present opportunities for future improvement.
In future work, we aim to optimize retrieval accuracy, refine emotional analysis, and expand the framework’s applicability to a broader range of LLM-generated stories to further validate its adaptability and generalizability.
Appendix A Implementation Settings
Table 3 to 6 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article. In the GPT playground, the full episode script is uploaded, and the prompt "Please evaluate episode n" is used to generate baseline results. When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode. As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
Appendix B Theoretical Analysis
B-A Proof of the Continuity Analysis of Critical Item State Correction
Assume a state space $\mathcal{M}$ containing states such as “active” and “destroyed” (with “lost” similarly defined). The system satisfies the Markov property:
$\begin{split}&P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\bigr{)}\\ &=P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\bigr{)}\,.\end{split}$ (1)
Denoting destroyed and lost as absorbing states, This mathematically means:
$P\big{(}S(t_{k})=\text{destroyed}\mid S(t_{k1})=\text{destroyed}\big{)}=1.$ (2)
Thus, no transition from destroyed or lost back to active or any non-terminal state is allowed.
We define the narrative entropy function of a state sequence by [10]. Let
$E(S_{0},S_{1},\dots,S_{T})=\sum_{k=0}^{T-1}\Delta E\big{(}S_{k},S_{k+1}\big{)},$ (3)
be the total narrative entropy over the sequence, where $\Delta E(S_{k},S_{k+1})$ represents the increase in entropy when transitioning from state $S_{k}$ to state $S_{k+1}$. It has $\Delta E\big{(}S_{k},S_{k+1}\big{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}$.
Then, the total narrative entropy over the sequence $\{S_{0},S_{1},\dots,S_{T}\}$ is:
$E(S_{0},S_{1},\dots,S_{T})=-\sum_{k=0}^{T-1}\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}.$ (4)
Because narratives are expected to be coherent, the theory requires that the entropy increases are minimal (ideally zero) along the legal state path. That is, for all legal transitions $S_{k}\to S_{k+1}$, we have
$\Delta E(S_{k},S_{k+1})=0.$ (5)
Now suppose, by way of contradiction, that an illegal transition is allowed. That is, there exists some $k$ for which
$P\big{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\big{)}>0.$ (6)
This transition introduces a jump in the narrative sequence. Let the entropy increase associated with this illegal transition be $\Delta E\big{(}\text{destroyed},\text{active}\big{)}$. Then we have:
$\begin{split}\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}&=\\ -\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\bigr{)}\Bigr{)}&>0\,.\end{split}$ (7)
Thus, the total narrative entropy for the sequence becomes
$\begin{split}E&=\sum_{j\neq k}\Delta E\bigl{(}S_{j},S_{j+1}\bigr{)}+\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}\\ &>0\,.\end{split}$ (8)
B-B Theory for complex question performance gain
B-B1 Preliminaries
Let $X$ be the original text (story), and $Y$ be the key information (summary or state representation), and $C$ be the relevant context for answering questions. The mutual information $I(X;Y)$ quantifies how much information $Y$ retains from $X$. Minimizing $I(X;Y)$ corresponds to compressing $X$ into $Y$. The mutual information $I(Y;C)$ quantifies how much information $Y$ shares with the relevant context $C$; maximizing this term ensures that crucial contextual elements are preserved.
According to the Information Bottleneck (IB) principle, when compressing $X$ to $Y$ while preserving information about $C$, one formulates the following objective:
$\min_{p(y|x)}I(X;Y)-\beta I(Y;C),$ (9)
where $\beta>0$ is a Lagrange multiplier balancing compression and preservation of relevant information.
In the SCORE system, The retrieval mechanism maximizes $I(Y;C)$; that is, it selects or emphasizes key information $Y$ which has high mutual information with the relevant context $C$. At the same time, by minimizing $I(X;Y)$, SCORE compresses the original text $X$ so that only the crucial parts (needed for QA) are retained.
Lemma B.1.
Suppose we want to learn a conditional distribution $p(y|x)$ so that the following objective $I(X;Y)-\beta I(Y;C)$ is minimized, then:
$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (10)
with
$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (11)
Proof.
We want to learn a conditional distribution $p(y|x)$ so that the following objective is minimized:
$J\left[p(y|x)\right]=I(X;Y)-\beta I(Y;C),\quad\beta>0,$ (12)
where
$I(X;Y)$ $=\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)},$ (13)
$I(Y;C)$ $=\sum_{y,c}p(y,c)\ln\frac{p(c|y)}{p(c)}.$
We use a Lagrange multiplier $\lambda(x)$ for each $x$ so that the normalization constraint $\sum_{y}p(y|x)=1$ holds. The Lagrangian functional defines as:
$\begin{split}\mathcal{L}\bigl{[}p(y\mid x)\bigr{]}&=\sum_{x,y}p(x)\,p(y\mid x)\,\ln\frac{p(y\mid x)}{p(y)}\\ &\quad-\beta\sum_{x,y,c}p(x)\,p(y\mid x)\,p(c\mid x)\,\ln\frac{p(c\mid y)}{p(c)}\\ &\quad+\sum_{x}\lambda(x)\Bigl{(}\sum_{y}p(y\mid x)-1\Bigr{)}\,.\end{split}$ (14)
Here $p(y)$ is the marginal given by $p(y)=\sum_{x}p(x)p(y|x)$, and similarly $p(y,c)=\sum_{x}p(x)p(c|x)p(y|x)$.
Differentiate $\mathcal{L}$ with respect to $p(y|x)$. For convenience, we note that the derivative of the first term with respect to $p(y|x)$ involves
$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}p(x)\,p(y\mid x)\,\ln\tfrac{p(y\mid x)}{p(y)}\Bigr{)}\\ &\quad=p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}\,.\end{split}$ (15)
where $p(y)$ itself depends on $p(y|x)$; however, in the variational calculation (assuming the standard IB procedure) those dependencies are handled self-consistently. Similarly, for the second term we have
$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}-\beta\,p(x)\,p(y\mid x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\Bigr{)}\\ &\quad=-\beta\,p(x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\,.\end{split}$ (16)
$\begin{split}&p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}-\beta\,p(x)\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\\ &\quad+\lambda(x)=0\,.\end{split}$ (17)
Dividing by $p(x)$ and we get:
$\ln p(y|x)=\ln p(y)+\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}-1-\frac{\lambda(x)}{p(x)}.$ (18)
Exponentiating both sides, we have:
$p(y|x)=p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right)\cdot\frac{1}{Z(x)},$ (19)
where we have defined the normalization factor $Z(x)=\exp\left(1+\frac{\lambda(x)}{p(x)}\right)$ such that $\sum_{y}p(y|x)=1$.
Thus, the optimal form is
$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (20)
with
$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (21)
Theorem B.2 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
Let $e_{c}$ denote the current episode and $e_{p}$ denote a past episode, both represented as vectors in an embedding space. Define the cosine similarity $S(e_{c},e_{p})$ and the sentiment score $\sigma(e)$ as given. Construct the retrieval kernel
$K(e_{c},e_{p})=\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right),$ (22)
and the optimal retrieval probability is
$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (23)
with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
The details and underlying proof are provided in the section B-B.
Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation. The focus is on narrative aspects such as character consistency, plot progression, emotional authenticity, and key item continuity. This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
B-B2 proof of theorem B.2
Proof.
Then, under the assumption that the key contextual information is reflected in both $S(e_{c},e_{p})$ and the sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$, the retrieval probability $p(e_{p}|e_{c})$ defined above is the optimal solution that minimizes the IB objective, i.e.,
$\min_{p(e_{p}|e_{c})}I(e_{c};e_{p})-\beta I(e_{p};C),$ (24)
subject to the normalization constraint $\sum_{e_{p}}p(e_{p}|e_{c})=1$.
From lemma B.1, by constructing the Lagrangian
$\begin{split}\mathcal{L}\bigl{[}p(e_{p}\mid e_{c})\bigr{]}&=\sum_{e_{c},e_{p}}p(e_{c})\,p(e_{p}\mid e_{c})\,\ln\frac{p(e_{p}\mid e_{c})}{p(e_{p})}\\ &\quad-\beta\,I(e_{p};C)\\ &\quad+\sum_{e_{c}}\lambda(e_{c})\Bigl{(}\sum_{e_{p}}p(e_{p}\mid e_{c})-1\Bigr{)}\,.\end{split}$ (25)
and taking the functional derivative with respect to $p(e_{p}|e_{c})$, one obtains the stationarity condition
$\ln p(e_{p}|e_{c})=\ln p(e_{p})+\beta\Delta(e_{c},e_{p})-1-\frac{\lambda(e_{c})}{p(e_{c})},$ (26)
where $\Delta(e_{c},e_{p})$ represents the contribution from the relevance term $I(e_{p};C)$.
Exponentiating and normalizing leads to
$p(e_{p}|e_{c})=\frac{p(e_{p})}{Z(e_{c})}\exp\left(\beta\Delta(e_{c},e_{p})\right).$ (27)
By identifying
$\beta\Delta(e_{c},e_{p})=\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau},$ (28)
and assuming a flat prior $p(e_{p})$ that can be absorbed into the normalizer $Z(e_{c})$, the expression reduces exactly to
$p(e_{p}|e_{c})=\frac{\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right)}{\sum_{e\in\mathcal{E}}\exp\left(\frac{S(e_{c},e)-\gamma\left|\sigma(e_{c})-\sigma(e)\right|}{\tau}\right)}.$ (29)
This concludes that the retrieval mechanism based on $K(e_{c},e_{p})$ optimally balances compression (minimizing $I(e_{c};e_{p})$) and context preservation (maximizing $I(e_{p};C)$), thereby minimizing the IB objective. ∎
B-B3 conclusion
In this setting, Minimizing the term $I(X;Y)=I(e_{c};e_{p})$ ensures that the key information $e_{p}$ is a compressed summary of the original data $e_{c}$, discarding redundant or uninformative parts. Maximizing $I(Y;C)$ ensures that the retrieved episodes $e_{p}$ retain the critical context $C$ (here, the narrative elements determined by both semantic similarity $S(e_{c},e_{p})$ and sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$).
The kernel $K(e_{c},e_{p})$ naturally balances these two goals through its exponential form, with tuning parameters $\tau$ and $\gamma$ controlling the sharpness of selection and the relative penalty on sentiment mismatch, respectively.
TABLE II: Evaluation Metrics Results
Evaluation Method BLEU-4
GPT4o with framework Upload Original Article files directly to GPT-4o playground
GPT4o playground Call the GPT-4o API, upload our processed files, and use our framework
GPT4o-mini with framework Upload files directly to GPT-4o-mini playground
GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
TABLE III: Evaluation of Episode 26 by GPT4o and use our prompt
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Evaluation of Current Episode (26):**
1. **Character Consistency:**
- **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative. His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature.
- **Wakana:** Her reactions are mostly consistent with her character, showing naivety and a desire to see the best in others. However, there is a discrepancy with the musical note hairpin-she fidgets with it despite it being lost in previous episodes, which is a continuity error.
- **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend.
2. **Plot Progression:**
- This episode contributes to the story by escalating the tension between the School President BF and the other characters. The suggestion of living together pushes the narrative of the BFs increasing control over Wakana. Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments.
3. **Continuity and Consistency in Story Elements:**
- There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost. This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil. Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition.
**Score: 3 (Fair)**
**Justification:**
The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters. However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence. These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
TABLE IV: Evaluation of Episode 26 in GPT4o playground
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Character Consistency:
In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits. The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature. His dialogue remains calculating and laced with undertones of jealousy, particularly when interacting with Kenji. The School President BFs obsessive behavior and desire to maintain control over Wakana are still central to his character, and his actions reflect this consistent personality trait.
Wakana maintains her characteristic obliviousness to the undercurrents of jealousy, which is also in line with her established naivety. However, theres a subtle evolution in her character as she shows more emotional depth and awareness of her surroundings, suggesting growth. Kenji, as always, remains timid and caught between his emotions for Wakana and fear of School President BF.
The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline.
Plot Progression:
This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji. The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school. The introduction of Kenjis potential transfer provides a dramatic twist, which escalates the conflict and sets up further developments.
The episode effectively builds suspense, especially with the School President BFs plans to force Kenji to leave, creating anticipation for future episodes. The plot thread involving Wakanas emotional conflict adds another layer of complexity to the narrative, keeping the audience engaged. The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines.
Continuity and Consistency in Story Elements:
There are no major continuity errors in this episode. The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts. The ongoing conflict between the School President BF and Kenji, as well as Wakanas central role, remains intact without any glaring contradictions or inconsistencies.
One minor point that could have been addressed more clearly is Wakanas hairpin, which has been a symbolic accessory throughout the series. While it appears briefly, its importance to the emotional tone of the episode could have been emphasized further, as it has been a key item in previous episodes. However, this does not significantly impact the narrative coherence.
Score: 4 (Good)
The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues. The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged. There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
TABLE V: Evaluation of Episode 26 by GPT4o-mini and use our prompt
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Critical Evaluation of Episode 26:**
1. **Character Consistency:**
- The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes. His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict.
- Wakanas innocence and naivety are also consistent, as she remains oblivious to the potential manipulation from the School President BF. However, there is an inconsistency with the mention of her musical note hairpin since it was previously noted as lost on a cliff. This could be an error unless it symbolizes her longing for the past.
2. **Plot Progression:**
- The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
- Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense.
- The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him.
3. **Continuity and Consistency in Story Elements:**
- The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost. This inconsistency undermines the narrative coherence.
- The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully.
**Score: 3 (Fair)**
**Justification:**
The episode is generally coherent with consistent character portrayals and logical plot progression. However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity. Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
TABLE VI: Evaluation of Episode 26 in GPT4o-mini playground
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Evaluation of Episode 26
Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal. His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes. Wakana, similarly, remains consistent in her obliviousness to the deeper conflict, maintaining her trusting and carefree personality, especially towards Kenji. There are no major inconsistencies in their behavior, and the progression of their personalities fits with the established dynamics of the story. However, Wakanas deep emotional response towards Kenji in this episode does suggest some growth or realization of his importance to her, which could potentially signal a shift in her character arc.
Plot Progression: This episode significantly advances the plot by heightening the tension between the School President BF and Kenji. The stakes are raised as the President overtly threatens Kenji, marking a clear escalation in his possessiveness and willingness to act on his jealousy. Kenjis reluctance to involve Wakana, despite being under intense pressure, adds complexity to the narrative, and the revelation that Kenji may be forced to leave the school contributes to the overall conflict. Wakanas emotional response hints at potential future conflict between her and the School President, as she may begin to question his actions and motivations. This scene effectively moves the story forward, setting the stage for a climax in their triangular relationship.
Continuity and Consistency in Story Elements: The episode maintains strong continuity with prior episodes, especially regarding key objects like Wakanas musical note hairpin, which has been a recurring symbol of her character. There are no glaring continuity errors, such as items reappearing without explanation. However, the escalating threat from the School President towards Kenji is beginning to stretch the believability of their relationship with Wakana, as her ongoing obliviousness to the intensity of the conflict between the two feels slightly contrived at this point. Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence.
Score: 4 (Good)
This episode is well-executed, with strong character consistency and meaningful plot advancement. The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances. There is also room for more explicit explanation of some story elements, such as the Presidents use of surveillance, to avoid straining the narratives credibility.

### v5

@@
 Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
-**footnotetext: Equal contribution.
+**footnotetext: Equal contribution.footnotetext: Corresponding author.
 I Introduction
@@
 with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
-The details and underlying proof are provided in the section B-B.
+The details and underlying proof are provided in the section˜B-B.
 Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation.
@@
 subject to the normalization constraint $\sum_{e_{p}}p(e_{p}|e_{c})=1$.
-From lemma B.1, by constructing the Lagrangian
+From lemma˜B.1, by constructing the Lagrangian
 $\begin{split}\mathcal{L}\bigl{[}p(e_{p}\mid e_{c})\bigr{]}&=\sum_{e_{c},e_{p}}p(e_{c})\,p(e_{p}\mid e_{c})\,\ln\frac{p(e_{p}\mid e_{c})}{p(e_{p})}\\ &\quad-\beta\,I(e_{p};C)\\ &\quad+\sum_{e_{c}}\lambda(e_{c})\Bigl{(}\sum_{e_{p}}p(e_{p}\mid e_{c})-1\Bigr{)}\,.\end{split}$ (25)
@@
 TABLE III: Evaluation of Episode 26 by GPT4o and use our prompt
-{mdframed}
-[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
-Score: 3
-Evaluation Reasoning:
-**Evaluation of Current Episode (26):**
-1. **Character Consistency:**
-- **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative.
+Score: 3 Evaluation Reasoning: **Evaluation of Current Episode (26):** 1.
+**Character Consistency:** - **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative.
 His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature.
@@
 - **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend.
-2. **Plot Progression:**
-- This episode contributes to the story by escalating the tension between the School President BF and the other characters.
+2.
+**Plot Progression:** - This episode contributes to the story by escalating the tension between the School President BF and the other characters.
 The suggestion of living together pushes the narrative of the BFs increasing control over Wakana.
 Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments.
-3. **Continuity and Consistency in Story Elements:**
-- There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost.
+3.
+**Continuity and Consistency in Story Elements:** - There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost.
 This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil.
 Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition.
-**Score: 3 (Fair)**
-**Justification:**
-The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters.
+**Score: 3 (Fair)** **Justification:** The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters.
 However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence.
@@
 TABLE IV: Evaluation of Episode 26 in GPT4o playground
-{mdframed}
-[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
-Character Consistency:
-In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits.
+Character Consistency: In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits.
 The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature.
@@
 The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline.
-Plot Progression:
-This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji.
+Plot Progression: This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji.
 The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school.
@@
 The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines.
-Continuity and Consistency in Story Elements:
-There are no major continuity errors in this episode.
+Continuity and Consistency in Story Elements: There are no major continuity errors in this episode.
 The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts.
@@
 However, this does not significantly impact the narrative coherence.
-Score: 4 (Good)
-The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues.
+Score: 4 (Good) The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues.
 The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged.
@@
 TABLE V: Evaluation of Episode 26 by GPT4o-mini and use our prompt
-{mdframed}
-[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
-Score: 3
-Evaluation Reasoning:
-**Critical Evaluation of Episode 26:**
-1. **Character Consistency:**
-- The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes.
+Score: 3 Evaluation Reasoning: **Critical Evaluation of Episode 26:** 1.
+**Character Consistency:** - The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes.
 His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict.
@@
 This could be an error unless it symbolizes her longing for the past.
-2. **Plot Progression:**
-- The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
+2.
+**Plot Progression:** - The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
 - Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense.
 - The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him.
-3. **Continuity and Consistency in Story Elements:**
-- The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost. This inconsistency undermines the narrative coherence.
+3.
+**Continuity and Consistency in Story Elements:** - The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost.
+This inconsistency undermines the narrative coherence.
 - The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully.
-**Score: 3 (Fair)**
-**Justification:**
-The episode is generally coherent with consistent character portrayals and logical plot progression.
+**Score: 3 (Fair)** **Justification:** The episode is generally coherent with consistent character portrayals and logical plot progression.
 However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity.
@@
 TABLE VI: Evaluation of Episode 26 in GPT4o-mini playground
-{mdframed}
-[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
-Evaluation of Episode 26
-Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal.
+Evaluation of Episode 26 Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal.
 His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes.
@@
 Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence.
-Score: 4 (Good)
-This episode is well-executed, with strong character consistency and meaningful plot advancement.
+Score: 4 (Good) This episode is well-executed, with strong character consistency and meaningful plot advancement.
 The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances.

### v6

@@
 In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies.
-By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF [1] and cosine similarity [2] to identify related episodes and enhance the overall story structure.
-Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
-Index Terms:
-Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
-**footnotetext: Equal contribution.footnotetext: Corresponding author.
-I Introduction
+By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach to identify related episodes and enhance the overall story structure.
+Experimental results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
+Index Terms—  Large language models, AI narrative, story structure.
+1 Introduction
 Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [1, 2, 3, 4, 5, 6].
-Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [9].
+Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [7, 8, 9].
 However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10].
 For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story.
-In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
+In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure [12].
 Similarly, [13] highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement.
 These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
+Fig.
+1: Overview of SCORE: (a) Extracts key item statuses per episode.
+(b) Conducts detailed analysis and summaries of each episode.
+(c) Uses RAG to answer user queries and resolve narrative inconsistencies.
 In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents.
@@
   • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
-  • We outperform baseline GPT model [16] in detecting continuity errors and maintaining overall narrative coherence.
-II Method
+Specially, we outperform baseline GPT model [17] in detecting continuity errors and maintaining overall narrative coherence.
+2 Method
 Our proposed method, SCORE, consists of three main components: (1) an LLM-based evaluation framework to assess the coherence of key story elements, (2) automatic generation of episode summaries to track plot development, and (3) a retrieval-augmented generation (RAG) approach that integrates the first two components, enabling enhanced user interaction and ensuring narrative consistency.
 As the framework is intended solely for academic research purposes, its use is consistent with the original access conditions of all incorporated tools and data sources.
-Figure 1: The SCORE framework for improving AI-generated story coherence.
-(a) Extracts key item statuses per episode.
-(b) Conducts detailed analysis and summaries of each episode.
-(c) Uses RAG to answer user queries and resolve narrative inconsistencies.
-II-A Continuity Analysis and Key Item Status Correction
+2.1 Continuity Analysis and Key Item Status Correction
 By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without narrative explanation.
@@
 This approach systematically corrects discrepancies in item states, ensuring that narrative continuity is preserved by preventing erroneous state transitions.
-II-B Key Item Interaction Analysis
+2.2 Key Item Interaction Analysis
 For each episode, we conduct a thorough evaluation by summarizing key plot points, character actions, and tracking interactions with important items.
@@
 The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
-II-C Similarity-Based Episode Evaluation and Sentiment Analysis
+Table 1: Performance Comparison of LLMs with/without SCORE Framework
+Model Consistency Coherence Item Status Complex QA
+GPT-4 83.21 84.32 0 82.34
+SCORE 85.61 ($\uparrow$2.4) 86.9 ($\uparrow$2.58) 98 ($\uparrow$98) 89.45 ($\uparrow$7.11)
+GPT-4o 86.78 82.21 0 76.32
+SCORE 88.68 ($\uparrow$1.9) 89.91 ($\uparrow$7.7) 96 ($\uparrow$96) 88.75 ($\uparrow$12.43)
+Claude3 84.6 80.9 0 69.45
+SCORE 87.2 ($\uparrow$2.6) 85.7 ($\uparrow$4.8) 93.1 ($\uparrow$93.1) 83.9 ($\uparrow$14.45)
+Gemini-Pro 82.2 83.4 0 78.40
+SCORE 85.2 ($\uparrow$3.2) 86.0 ($\uparrow$3.0) 95.0 ($\uparrow$95.0) 85.5 ($\uparrow$7.10)
+Qwen-14B 79.3 76.2 0 25.15
+SCORE 84.5 ($\uparrow$5.2) 79.1 ($\uparrow$2.9) 86.3 ($\uparrow$86.3) 61.2 ($\uparrow$36.05)
+GPT-4o-mini 77.5 75.8 0 23.40
+SCORE 81.9 ($\uparrow$4.4) 76.6 ($\uparrow$0.8) 79.8 ($\uparrow$79.8) 60.1 ($\uparrow$36.7)
+Llama-13B 71.3 69.8 0 18.72
+SCORE 79.1 ($\uparrow$7.8) 73.4 ($\uparrow$3.6) 76.2 ($\uparrow$76.2) 57.3 ($\uparrow$38.58)
+2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
 We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries.
@@
 This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
-For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores between the current episode or query and all other episodes in the vector space.
+For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores [19] between the current episode or query and all other episodes in the vector space.
 Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$.
@@
 This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
-TABLE I: Performance comparison of models with and without SCORE
-Model Consistency Coherence Item Status Complex Question
-GPT-4o-mini 78.2 76.7 0 24.56
-+ SCORE 82.6 (↑4.4) 77.5 (↑0.8) 80.5 (↑80.5) 63.0 (↑38.44)
-GPT-4o 86.78 82.21 0 76.32
-+ SCORE 88.68 (↑1.9) 89.91 (↑7.7) 96 (↑96) 88.75 (↑12)
-GPT-4 83.21 84.32 0 82.34
-+ SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
-III Experiments
+3 Experiments
 To evaluate the effectiveness of SCORE, we conducted experiments on stories generated by large language models (LLMs).
 These experiments assessed the framework’s ability to maintain narrative coherence, detect continuity errors, and ensure emotional consistency throughout episodic storytelling.
-III-A Dataset Preparation
-We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy.
-Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements.
-This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
-III-B Baselines
+Table 2: Ablation results for our complete configuration (Full SCORE) versus removing each module.
+Configuration Consistency Coherence Item Acc. Complex QA
+Full SCORE 89.7 91.2 98.3 88.5
+w/o Dynamic Tracking 72.1 ($\downarrow$17.6) 85.4 ($\downarrow$5.8) 61.2 ($\downarrow$37.1) 67.3 ($\downarrow$21.2)
+w/o Context Summary 83.2 ($\downarrow$6.5) 68.7 ($\downarrow$22.5) 89.1 ($\downarrow$9.2) 71.4 ($\downarrow$17.1)
+w/o Hybrid Retrieval 86.4 ($\downarrow$3.3) 88.9 ($\downarrow$2.3) 94.2 ($\downarrow$4.1) 77.6 ($\downarrow$10.9)
+w/o Sentiment 87.1 ($\downarrow$2.6) 89.5 ($\downarrow$1.7) 96.8 ($\downarrow$1.5) 82.3 ($\downarrow$6.2)
+Fig. 2: Case study.
+3.1 Dataset Preparation
+We constructed a story dataset of 5,000 episodes from 1,000 GPT-generated stories evenly distributed across four genres (science fiction, drama, fantasy, comedy).
+Each story contains 10–15 episodes (avg.
+12, $\sim$2,000 tokens), generated using GPT-3.5 and GPT-4 outputs with 15 diverse prompt templates per genre.
+Quality was ensured via manual filtering by three annotators ($\kappa$=0.78), and complexity stratification included multi-character interactions (30%), non-linear timelines (20%), and symbolic systems (50%).
+For broader validation, SCORE was also tested on NarrativeQA, BookCorpus, and WP-STORIES, with genre and episode length distributions standardized for cross-dataset comparison.
+This hybrid setup balances synthetic and human-authored narratives, ensuring robustness across industrial and academic benchmarks.
+3.2 Metrics
+We evaluated the framework using several key metrics: narrative coherence, assessed by examining logical consistency in character behavior and plot development to avoid continuity errors; consistency, defined as the fraction of responses not conflicting with prior information; coherence, measured as the average semantic or logical flow score across responses; item status, calculated as the proportion of required elements present in each response; and complex question, defined as the fraction of correctly answered complex questions.
+Finally, we compared the evaluation scores from different methods to measure the stability of the framework.
+3.3 Evaluation Process
+For each episode, evaluation proceeded in two stages.
+First, we directly uploaded files to ChatGPT (GPT-4o-mini, GPT-4o, GPT-4) to test baseline narrative assessment without prompts.
+Second, we performed detailed evaluation using preprocessed files with RAG support, where FAISS-based cosine similarity retrieved semantically and emotionally aligned episodes to provide context.
+This enriched context was then incorporated into GPT prompts for more accurate key-item tracking and narrative evaluation.
+3.4 Baselines
 We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini.
@@
 For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
-III-C Evaluation Process
-For each episode in the dataset, the evaluation process involved several key steps.
-First, we conducted an initial evaluation by directly uploading the files to ChatGPT, testing if GPT could assess the story correctly without any prompts.
-We used GPT-4o-mini, GPT-4o, and GPT-4 for narrative evaluation.
-Next, we configured the LLM for a more detailed evaluation, using our preprocessed files to ensure accurate tracking of key items across episodes.
-We also employed the Retrieval-Augmented Generation (RAG) framework to retrieve episodes similar to the one being evaluated.
-By using FAISS to calculate cosine similarity scores between episodes, the system was able to retrieve episodes that were semantically and emotionally aligned with the current one, providing additional context.
-This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
-III-D Metrics
-We evaluated the framework based on several key metrics.
-Narrative coherence was assessed by examining how well the framework maintains the logic of the story.
-Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected.
-Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
-III-E Results
-Our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies.
+3.5 Main Results
+As shown in Table 1, our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies.
 Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression.
@@
 When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
-IV Limitations
-While our SCORE framework offers significant improvements, it also presents several limitations:
-  • Dependence on Retrieval Accuracy: The effectiveness of SCORE relies on accurately retrieving relevant episodes.
-Errors in similarity calculation or sentiment analysis may lead to the exclusion of important contextual information, affecting the overall evaluation quality.
-  • Challenges in Capturing Emotional Nuances: Sentiment analysis may not fully capture the complexity of emotions, potentially missing subtle shifts or deeper emotional layers within the narrative.
-  • Resource Intensive: Despite optimizations, SCORE still requires considerable computational resources for similarity calculations and sentiment analysis, which can be limiting for large datasets.
-  • Generalizability: SCORE has been primarily tested on GPT-4-generated stories, and its performance on narratives from other models or with different structures remains an open question.
-V Conclusion
-We introduced SCORE, a novel LLM-based evaluation framework aimed at improving the coherence and emotional consistency of AI-generated stories.
-By incorporating Retrieval-Augmented Generation (RAG), TF-IDF, cosine similarity, and sentiment analysis, our method retrieves contextually relevant information, enabling more accurate and detailed assessments.
-Experimental results demonstrate that SCORE outperforms traditional methods by effectively identifying continuity errors, maintaining narrative coherence, and providing valuable insights for episodic storytelling.
-However, limitations such as dependency on retrieval accuracy and computational demands present opportunities for future improvement.
-In future work, we aim to optimize retrieval accuracy, refine emotional analysis, and expand the framework’s applicability to a broader range of LLM-generated stories to further validate its adaptability and generalizability.
-Appendix A Implementation Settings
-Table 3 to 6 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article.
-In the GPT playground, the full episode script is uploaded, and the prompt "Please evaluate episode n" is used to generate baseline results.
-When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode.
-As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
-Appendix B Theoretical Analysis
-B-A Proof of the Continuity Analysis of Critical Item State Correction
-Assume a state space $\mathcal{M}$ containing states such as “active” and “destroyed” (with “lost” similarly defined). The system satisfies the Markov property:
-$\begin{split}&P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\bigr{)}\\ &=P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\bigr{)}\,.\end{split}$ (1)
-Denoting destroyed and lost as absorbing states, This mathematically means:
-$P\big{(}S(t_{k})=\text{destroyed}\mid S(t_{k1})=\text{destroyed}\big{)}=1.$ (2)
-Thus, no transition from destroyed or lost back to active or any non-terminal state is allowed.
-We define the narrative entropy function of a state sequence by [10]. Let
-$E(S_{0},S_{1},\dots,S_{T})=\sum_{k=0}^{T-1}\Delta E\big{(}S_{k},S_{k+1}\big{)},$ (3)
-be the total narrative entropy over the sequence, where $\Delta E(S_{k},S_{k+1})$ represents the increase in entropy when transitioning from state $S_{k}$ to state $S_{k+1}$.
-It has $\Delta E\big{(}S_{k},S_{k+1}\big{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}$.
-Then, the total narrative entropy over the sequence $\{S_{0},S_{1},\dots,S_{T}\}$ is:
-$E(S_{0},S_{1},\dots,S_{T})=-\sum_{k=0}^{T-1}\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}.$ (4)
-Because narratives are expected to be coherent, the theory requires that the entropy increases are minimal (ideally zero) along the legal state path.
-That is, for all legal transitions $S_{k}\to S_{k+1}$, we have
-$\Delta E(S_{k},S_{k+1})=0.$ (5)
-Now suppose, by way of contradiction, that an illegal transition is allowed. That is, there exists some $k$ for which
-$P\big{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\big{)}>0.$ (6)
-This transition introduces a jump in the narrative sequence. Let the entropy increase associated with this illegal transition be $\Delta E\big{(}\text{destroyed},\text{active}\big{)}$. Then we have:
-$\begin{split}\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}&=\\ -\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\bigr{)}\Bigr{)}&>0\,.\end{split}$ (7)
-Thus, the total narrative entropy for the sequence becomes
-$\begin{split}E&=\sum_{j\neq k}\Delta E\bigl{(}S_{j},S_{j+1}\bigr{)}+\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}\\ &>0\,.\end{split}$ (8)
-B-B Theory for complex question performance gain
-B-B1 Preliminaries
-Let $X$ be the original text (story), and $Y$ be the key information (summary or state representation), and $C$ be the relevant context for answering questions.
-The mutual information $I(X;Y)$ quantifies how much information $Y$ retains from $X$.
-Minimizing $I(X;Y)$ corresponds to compressing $X$ into $Y$.
-The mutual information $I(Y;C)$ quantifies how much information $Y$ shares with the relevant context $C$; maximizing this term ensures that crucial contextual elements are preserved.
-According to the Information Bottleneck (IB) principle, when compressing $X$ to $Y$ while preserving information about $C$, one formulates the following objective:
-$\min_{p(y|x)}I(X;Y)-\beta I(Y;C),$ (9)
-where $\beta>0$ is a Lagrange multiplier balancing compression and preservation of relevant information.
-In the SCORE system, The retrieval mechanism maximizes $I(Y;C)$; that is, it selects or emphasizes key information $Y$ which has high mutual information with the relevant context $C$.
-At the same time, by minimizing $I(X;Y)$, SCORE compresses the original text $X$ so that only the crucial parts (needed for QA) are retained.
-Lemma B.1.
-Suppose we want to learn a conditional distribution $p(y|x)$ so that the following objective $I(X;Y)-\beta I(Y;C)$ is minimized, then:
-$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (10)
-with
-$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (11)
-Proof.
-We want to learn a conditional distribution $p(y|x)$ so that the following objective is minimized:
-$J\left[p(y|x)\right]=I(X;Y)-\beta I(Y;C),\quad\beta>0,$ (12)
-where
-$I(X;Y)$ $=\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)},$ (13)
-$I(Y;C)$ $=\sum_{y,c}p(y,c)\ln\frac{p(c|y)}{p(c)}.$
-We use a Lagrange multiplier $\lambda(x)$ for each $x$ so that the normalization constraint $\sum_{y}p(y|x)=1$ holds. The Lagrangian functional defines as:
-$\begin{split}\mathcal{L}\bigl{[}p(y\mid x)\bigr{]}&=\sum_{x,y}p(x)\,p(y\mid x)\,\ln\frac{p(y\mid x)}{p(y)}\\ &\quad-\beta\sum_{x,y,c}p(x)\,p(y\mid x)\,p(c\mid x)\,\ln\frac{p(c\mid y)}{p(c)}\\ &\quad+\sum_{x}\lambda(x)\Bigl{(}\sum_{y}p(y\mid x)-1\Bigr{)}\,.\end{split}$ (14)
-Here $p(y)$ is the marginal given by $p(y)=\sum_{x}p(x)p(y|x)$, and similarly $p(y,c)=\sum_{x}p(x)p(c|x)p(y|x)$.
-Differentiate $\mathcal{L}$ with respect to $p(y|x)$. For convenience, we note that the derivative of the first term with respect to $p(y|x)$ involves
-$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}p(x)\,p(y\mid x)\,\ln\tfrac{p(y\mid x)}{p(y)}\Bigr{)}\\ &\quad=p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}\,.\end{split}$ (15)
-where $p(y)$ itself depends on $p(y|x)$; however, in the variational calculation (assuming the standard IB procedure) those dependencies are handled self-consistently.
-Similarly, for the second term we have
-$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}-\beta\,p(x)\,p(y\mid x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\Bigr{)}\\ &\quad=-\beta\,p(x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\,.\end{split}$ (16)
-$\begin{split}&p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}-\beta\,p(x)\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\\ &\quad+\lambda(x)=0\,.\end{split}$ (17)
-Dividing by $p(x)$ and we get:
-$\ln p(y|x)=\ln p(y)+\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}-1-\frac{\lambda(x)}{p(x)}.$ (18)
-Exponentiating both sides, we have:
-$p(y|x)=p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right)\cdot\frac{1}{Z(x)},$ (19)
-where we have defined the normalization factor $Z(x)=\exp\left(1+\frac{\lambda(x)}{p(x)}\right)$ such that $\sum_{y}p(y|x)=1$.
-Thus, the optimal form is
-$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (20)
-with
-$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (21)
-Theorem B.2 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
-Let $e_{c}$ denote the current episode and $e_{p}$ denote a past episode, both represented as vectors in an embedding space.
-Define the cosine similarity $S(e_{c},e_{p})$ and the sentiment score $\sigma(e)$ as given.
-Construct the retrieval kernel
-$K(e_{c},e_{p})=\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right),$ (22)
-and the optimal retrieval probability is
-$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (23)
-with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
-The details and underlying proof are provided in the section˜B-B.
-Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation.
-The focus is on narrative aspects such as character consistency, plot progression, emotional authenticity, and key item continuity.
-This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
-B-B2 proof of theorem B.2
-Proof.
-Then, under the assumption that the key contextual information is reflected in both $S(e_{c},e_{p})$ and the sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$, the retrieval probability $p(e_{p}|e_{c})$ defined above is the optimal solution that minimizes the IB objective, i.e.,
-$\min_{p(e_{p}|e_{c})}I(e_{c};e_{p})-\beta I(e_{p};C),$ (24)
-subject to the normalization constraint $\sum_{e_{p}}p(e_{p}|e_{c})=1$.
-From lemma˜B.1, by constructing the Lagrangian
-$\begin{split}\mathcal{L}\bigl{[}p(e_{p}\mid e_{c})\bigr{]}&=\sum_{e_{c},e_{p}}p(e_{c})\,p(e_{p}\mid e_{c})\,\ln\frac{p(e_{p}\mid e_{c})}{p(e_{p})}\\ &\quad-\beta\,I(e_{p};C)\\ &\quad+\sum_{e_{c}}\lambda(e_{c})\Bigl{(}\sum_{e_{p}}p(e_{p}\mid e_{c})-1\Bigr{)}\,.\end{split}$ (25)
-and taking the functional derivative with respect to $p(e_{p}|e_{c})$, one obtains the stationarity condition
-$\ln p(e_{p}|e_{c})=\ln p(e_{p})+\beta\Delta(e_{c},e_{p})-1-\frac{\lambda(e_{c})}{p(e_{c})},$ (26)
-where $\Delta(e_{c},e_{p})$ represents the contribution from the relevance term $I(e_{p};C)$.
-Exponentiating and normalizing leads to
-$p(e_{p}|e_{c})=\frac{p(e_{p})}{Z(e_{c})}\exp\left(\beta\Delta(e_{c},e_{p})\right).$ (27)
-By identifying
-$\beta\Delta(e_{c},e_{p})=\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau},$ (28)
-and assuming a flat prior $p(e_{p})$ that can be absorbed into the normalizer $Z(e_{c})$, the expression reduces exactly to
-$p(e_{p}|e_{c})=\frac{\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right)}{\sum_{e\in\mathcal{E}}\exp\left(\frac{S(e_{c},e)-\gamma\left|\sigma(e_{c})-\sigma(e)\right|}{\tau}\right)}.$ (29)
-This concludes that the retrieval mechanism based on $K(e_{c},e_{p})$ optimally balances compression (minimizing $I(e_{c};e_{p})$) and context preservation (maximizing $I(e_{p};C)$), thereby minimizing the IB objective.
-∎
-B-B3 conclusion
-In this setting, Minimizing the term $I(X;Y)=I(e_{c};e_{p})$ ensures that the key information $e_{p}$ is a compressed summary of the original data $e_{c}$, discarding redundant or uninformative parts.
-Maximizing $I(Y;C)$ ensures that the retrieved episodes $e_{p}$ retain the critical context $C$ (here, the narrative elements determined by both semantic similarity $S(e_{c},e_{p})$ and sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$).
-The kernel $K(e_{c},e_{p})$ naturally balances these two goals through its exponential form, with tuning parameters $\tau$ and $\gamma$ controlling the sharpness of selection and the relative penalty on sentiment mismatch, respectively.
-TABLE II: Evaluation Metrics Results
-Evaluation Method BLEU-4
-GPT4o with framework Upload Original Article files directly to GPT-4o playground
-GPT4o playground Call the GPT-4o API, upload our processed files, and use our framework
-GPT4o-mini with framework Upload files directly to GPT-4o-mini playground
-GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
-TABLE III: Evaluation of Episode 26 by GPT4o and use our prompt
-Score: 3 Evaluation Reasoning: **Evaluation of Current Episode (26):** 1.
-**Character Consistency:** - **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative.
-His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature.
-- **Wakana:** Her reactions are mostly consistent with her character, showing naivety and a desire to see the best in others.
-However, there is a discrepancy with the musical note hairpin-she fidgets with it despite it being lost in previous episodes, which is a continuity error.
-- **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend.
-2.
-**Plot Progression:** - This episode contributes to the story by escalating the tension between the School President BF and the other characters.
-The suggestion of living together pushes the narrative of the BFs increasing control over Wakana.
-Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments.
-3.
-**Continuity and Consistency in Story Elements:** - There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost.
-This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil.
-Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition.
-**Score: 3 (Fair)** **Justification:** The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters.
-However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence.
-These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
-TABLE IV: Evaluation of Episode 26 in GPT4o playground
-Character Consistency: In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits.
-The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature.
-His dialogue remains calculating and laced with undertones of jealousy, particularly when interacting with Kenji.
-The School President BFs obsessive behavior and desire to maintain control over Wakana are still central to his character, and his actions reflect this consistent personality trait.
-Wakana maintains her characteristic obliviousness to the undercurrents of jealousy, which is also in line with her established naivety.
-However, theres a subtle evolution in her character as she shows more emotional depth and awareness of her surroundings, suggesting growth.
-Kenji, as always, remains timid and caught between his emotions for Wakana and fear of School President BF.
-The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline.
-Plot Progression: This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji.
-The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school.
-The introduction of Kenjis potential transfer provides a dramatic twist, which escalates the conflict and sets up further developments.
-The episode effectively builds suspense, especially with the School President BFs plans to force Kenji to leave, creating anticipation for future episodes.
-The plot thread involving Wakanas emotional conflict adds another layer of complexity to the narrative, keeping the audience engaged.
-The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines.
-Continuity and Consistency in Story Elements: There are no major continuity errors in this episode.
-The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts.
-The ongoing conflict between the School President BF and Kenji, as well as Wakanas central role, remains intact without any glaring contradictions or inconsistencies.
-One minor point that could have been addressed more clearly is Wakanas hairpin, which has been a symbolic accessory throughout the series.
-While it appears briefly, its importance to the emotional tone of the episode could have been emphasized further, as it has been a key item in previous episodes.
-However, this does not significantly impact the narrative coherence.
-Score: 4 (Good) The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues.
-The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged.
-There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
-TABLE V: Evaluation of Episode 26 by GPT4o-mini and use our prompt
-Score: 3 Evaluation Reasoning: **Critical Evaluation of Episode 26:** 1.
-**Character Consistency:** - The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes.
-His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict.
-- Wakanas innocence and naivety are also consistent, as she remains oblivious to the potential manipulation from the School President BF.
-However, there is an inconsistency with the mention of her musical note hairpin since it was previously noted as lost on a cliff.
-This could be an error unless it symbolizes her longing for the past.
-2.
-**Plot Progression:** - The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
-- Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense.
-- The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him.
-3.
-**Continuity and Consistency in Story Elements:** - The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost.
-This inconsistency undermines the narrative coherence.
-- The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully.
-**Score: 3 (Fair)** **Justification:** The episode is generally coherent with consistent character portrayals and logical plot progression.
-However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity.
-Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
-TABLE VI: Evaluation of Episode 26 in GPT4o-mini playground
-Evaluation of Episode 26 Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal.
-His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes.
-Wakana, similarly, remains consistent in her obliviousness to the deeper conflict, maintaining her trusting and carefree personality, especially towards Kenji.
-There are no major inconsistencies in their behavior, and the progression of their personalities fits with the established dynamics of the story.
-However, Wakanas deep emotional response towards Kenji in this episode does suggest some growth or realization of his importance to her, which could potentially signal a shift in her character arc.
-Plot Progression: This episode significantly advances the plot by heightening the tension between the School President BF and Kenji.
-The stakes are raised as the President overtly threatens Kenji, marking a clear escalation in his possessiveness and willingness to act on his jealousy.
-Kenjis reluctance to involve Wakana, despite being under intense pressure, adds complexity to the narrative, and the revelation that Kenji may be forced to leave the school contributes to the overall conflict.
-Wakanas emotional response hints at potential future conflict between her and the School President, as she may begin to question his actions and motivations.
-This scene effectively moves the story forward, setting the stage for a climax in their triangular relationship.
-Continuity and Consistency in Story Elements: The episode maintains strong continuity with prior episodes, especially regarding key objects like Wakanas musical note hairpin, which has been a recurring symbol of her character.
-There are no glaring continuity errors, such as items reappearing without explanation.
-However, the escalating threat from the School President towards Kenji is beginning to stretch the believability of their relationship with Wakana, as her ongoing obliviousness to the intensity of the conflict between the two feels slightly contrived at this point.
-Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence.
-Score: 4 (Good) This episode is well-executed, with strong character consistency and meaningful plot advancement.
-The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances.
-There is also room for more explicit explanation of some story elements, such as the Presidents use of surveillance, to avoid straining the narratives credibility.
+3.6 Ablation Studies
+LLM type.
+Table 1 shows that SCORE improves performance across all model families, with especially large gains for open-source models.
+Consistency increases up to 7.8% (Llama-13B), coherence up to 7.7% (GPT-4o), and item status recognition approaches 98%.
+Open-source models (e.g., Qwen-14B, Llama-13B) see the largest QA gains (30–38 points), while commercial models (e.g., GPT-4, Claude 3) also benefit, though more modestly.
+Notably, SCORE boosts GPT-4o beyond GPT-4 in coherence, underscoring its scalability.
+Configuration analysis.
+Table 2 shows that Dynamic Tracking is critical, with its removal causing the largest drops in consistency (–17.6 points) and item accuracy (–37.1 points).
+Context Summary is similarly vital for coherence (–22.5%).
+Hybrid Retrieval and Sentiment contribute positively but with smaller impacts.
+3.7 Case Study
+We generated several dozen stories with SCORE, averaging 555 tokens and 13.2s runtime per episode.
+Evaluation across four metrics—character consistency, plot progression, emotional realism, and continuity—yielded average scores of 3–4/5 (Figure 2), indicating generally cohesive narratives.
+4 Conclusion
+We present SCORE, a novel LLM-based framework that enhances long-term coherence and emotional consistency in AI narratives.
+By combining Dynamic State Tracking, Context-Aware Summarization, and Hybrid Retrieval within a RAG pipeline, SCORE achieves substantial improvements in coherence, stability, and hallucination reduction across multi-genre datasets.
+Its modular design ensures scalability and multi-LLM compatibility, though challenges remain in retrieval accuracy and computational efficiency.